# Power Variable Projection for Initialization-Free Large-Scale Bundle Adjustment

### v1

@@ 4.3 Power Riemannian manifold optimization
As the projective refinement step in Equation 7 involves both camera matrices and 3D landmarks in homogeneous forms, we exhibit local scale freedom for both camera and landmark parameters. This necessitates incorporation of the Riemannian manifold optimization framework without which the linearized system of equations is always rank-deficient and unsolvable. While a complete theoretical overview of such framework can be found in [1], we formalize Equation 7 as:
@@
and then Equation 20 becomes:
$\begin{pmatrix}\tilde{U}_{\lambda}&&\tilde{W}\\ \tilde{W}^{\top}&&\tilde{V}_{\lambda}\end{pmatrix}\begin{pmatrix}\Delta x_{p}\\ \Delta x_{l}\end{pmatrix}=-\begin{pmatrix}\tilde{b}_{p}\\ \tilde{b}_{l}\end{pmatrix}\,,$ (22)
@@
$\tilde{U}_{\lambda}=\tilde{J}_{p}^{\top}\tilde{J}_{p}+D_{p}^{\top}\tilde{\lambda}_{p}D_{p}\,,$ (23)
$\tilde{V}_{\lambda}=\tilde{J}_{l}^{\top}\tilde{J}_{l}+D_{l}^{\top}\tilde{\lambda}_{l}D_{l}\,,$ (24)
@@
Figure 2: Performance profiles for all real-world BAL problems for solving the first stage Equation 5. Given a tolerance $\tau\in\{0.01,0.003,0.001\}$, it represents the percentage of solved problems ($y$-axis) with relative runtime $\alpha$ ($x$-axis). Expansion methods PoVar and PoBA are very competitive. Our solver PoVar is very competitive, and most notably for the highest accuracy $\tau=0.001$ and $\tau=0.003$.
Figure 3: Convergence plots of Dubrovnik-253 (left) from BAL datasets with 253 poses and Ladybug-412 with 412 poses, for solving the first stage (Equation 5). The dotted lines correspond to cost thresholds for tolerance $\tau\in\{0.01,0.003,0.001\}$.
Figure 4: Performance profiles for all real-world BAL problems for solving the first two stages Equation 5 and Equation 7. In each combination, the first solver is used to solve the first stage, and the second solver is used for the second stage. Our proposed combination PoVar followed by a Riemannian expansion method outperforms by a large margin compared to all its competitors. Let emphasize that our proposed Riemannian expansion method RiPoBA outperforms the iterative baseline in all cases, given a same solver for the first stage.
Figure 5: Convergence plots of Ladybug-460 (left) from BAL datasets with 460 poses and Venice-1672 with 1672 poses, for solving the second stage (Equation 7). The dotted lines correspond to cost thresholds for tolerance $\tau\in\{0.01,0.003,0.001\}$. Note that for fair comparison, the initial cost is derived before the first stage. The second cost – that may be higher than the initial one, is the initial cost of the second stage, after the first stage has been run. The runtime includes the time spent to solve the first stage.
@@ 5.1 Implementation
We implement pOSE [^1] We use our custom implementation for comparisons, as the implementation of [11] is not publicly available. [11], PoVar and RiPoBA framework in C++, directly on the publicly available implementation of PoBA  [^2] https://github.com/simonwebertum/poba [25]. That leads to fair comparisons with this recent and very challenging solver. pOSE differs from PoVar by the use of a direct sparse Cholesky factorization. We also compare to VarPro with the conjugate gradients algorithm, preconditioned with Schur-Jacobi preconditioner, called iterative in our experiments. For the second stage, we compare RiPoBA to the conjugate gradients algorithm preconditioned by Schur-Jacobi preconditioners with Riemannian manifold optimization framework, called RiPCG in our experiments. Except the solver itself, all implementations share much of the code with [25]. We run experiments on MacOS 14.2.1 with an Intel Core i5 and 4 cores at 2GHz.
@@ 5.2 Experimental settings
For each stage, we set the maximum number of iterations to $50$, stopping earlier if a relative function tolerance of $10^{-6}$ is achieved. The damping parameter $\lambda$ starts for each stage at $10^{-4}$ and is updated accordingly to the success or failure of the iteration. For expansion methods, we set the maximal order of power series to $20$ and a threshold to $0.01$. For iterative methods, we set the maximal number of inner iterations to $500$. We set the coefficient $\eta$ for the pOSE residuals (Equation 6) to $0.1$.
Figure 6: Example of dense landmark block in the tangent space, with three observations of the considered landmark. We project the initial dense landmark block (left figure) onto the tangent space by applying to each $i$-th pose Jacobian $J_{p_{i}}\in\mathbb{R}^{2\times 12}$ the projection $\tilde{x}_{p_{i}}^{\perp}$, and to each $j$-th landmark Jacobian $J_{l_{j}}\in\mathbb{R}^{2\times 4}$ the projection $\tilde{x}_{l_{j}}^{\perp}$. The resulting pose and landmark Jacobians in the tangent space (right figure) respectively belong to $\mathbb{R}^{2\times 11}$ and $\mathbb{R}^{2\times 3}$.
@@
As in [25], we leverage the special structure of BA problem and propose a memory-efficient storage. We design a dense projected landmark blocks. In particular, we apply on each row associated to a landmark the block-matrices of the projection $\tilde{x}^{\perp}_{l_{j}},\tilde{x}^{\perp}_{p_{i}}$ corresponding to the landmark and to the cameras stored in the considered dense landmark block (see Figure 6).
@@
We extensively evaluate our solver and the baselines on the $97$ real-world bundle adjustment problems from the BAL project page [3]. The number of poses goes from $16$ to $13682$. We refer the reader to Supplemental for further details about these problems. For each problem, we only keep the observation measurements. Pose parameters are randomly drawn from an isotropic Gaussian distribuaton with mean 00 and variance $1$, and landmark parameters are deduced with Equation 8. Notably and contrary to previous works on initialization-free BA, each solver is ran on the same randomized problem, for fair comparisons.
@@ 5.3 Performance profile
To present a broad and fair analysis, we jointly evaluate both runtime and accuracy with performance profiles [8]. Given a solver, the performance profile maps the relative runtime $\alpha$ to the percentage of problems solved with accuracy $\tau$. Graphically, the performance profile of a given solver is the percentage of problems solved faster than the relative runtime $\alpha$ on the $x$-axis. Let be $S$ and $P$ the sets of solvers and problems, respectively. In practice, we can define the objective threshold for a problem $p$ by:
@@ 5.4 Analysis
Figure 2 shows the performance profiles for all BAL datasets with tolerances $\tau\in\{0.01,0.003,0.001\}$ to solve Equation 5. As expected, the direct factorization solver (dashed green curve) used in Hong et al. [11] shows poor performance. Our solver PoVar (blue curve) challenges PoBA for the largest tolerance $0.01$, and is by far the most competitive solver for the smallest tolerance $\tau=0.001$, that is for the highest accuracy. For $\tau=0.003$, PoVar slightly outperforms PoBA. Let also highlight that our solver clearly outperforms the two competitors associated to the VarPro algorithm, iterative (dashed red) and direct factorization.
Figure 3 illustrates on two examples the cost decrease during the first stage. Notably, PoVar (blue curve) shows a much smoother convergence than its main challenger PoBA, that tends to stuck during the first iterations. By considering the intersections of the solvers with the cost thresholds (dashed grey lines), it also highlights the slowness of the popular iterative method compared to expansion methods, in terms of runtime.
@@
Figure 4 shows the performance profiles for all BAL datasets with tolerances $\tau\in\{0.01,0.003,0.001\}$ to solve the first two stages, that are Equation 5 followed by Equation 7. Note that we compare in this experiment the cost of the second stage only, as the first stage is only used to get an approximated initialization for the projective formulation. As direct factorization shows very poor performance during the first stage, we only take into account the most promising combinations of solvers. PoVar followed by RiPoBA (blue curve) clearly outperforms all the competitors with a large margin, both in terms of runtime and accuracy. The combinations with RiPoBA outperform RiPCG for the highest accuracy $\tau=0.001$ for all relative time greater than $\alpha=2$, that reflects the better convergence of Riemannian expansion method compared to RiPCG. Last but not least, given a same solver during first stage, RiPoBA clearly outperforms RiPCG during second stage, for all tolerances.
Figure 5 illustrates on two examples the cost decrease during the second stage. On the left figure, associated to Ladybug-460, the best two solvers in terms of final convergence are built on our framework RiPoBA. On the right figure, associated to Venice-1672, all the solvers using RiPoBA converge to a smaller error than their iterative competitor RiPCG. By considering the intersection between the solvers and the cost thresholds (dashed grey lines), the combination PoVar-RiPoBA clearly outperforms all the other combinations, and most notably for the largest problem with $1672$ poses.
@@
Overall, the experiments emphasize the high effectivity of our solvers PoVar and RiPoBA, during both first and second stages of the stratified BA problem. Concerning the first stage, the convergence of PoVar is much smoother than its competitors, that explains its larger speed with respect to PoBA to reach the cost thresholds, albeit both are built with power series. Regarding the second stage, for a same given solver in the first stage, our proposed RiPoBA clearly outperforms the popular preconditioned conjugate gradients with Riemannian manifold optimization framework in terms of speed and accuracy, especially when coupled with PoVar.
@@ 6 Conclusion
We introduce a novel approach to address the scalability challenge for initialization-free bundle adjustment. Our proposed Power Variable Projection (PoVar) algorithm, theoretically justified, offers new insights to this uncharted problem. By extending recent inverse expansion techniques to the VarPro algorithm on one hand, and to Riemannian manifold optimization on the other hand, we have demonstrated the capability to efficiently solve large-scale stratified BA problem with thousands of cameras. Notably, we achieve state-of-the-art results in terms of speed and accuracy on the real-world BAL dataset. While initialization-free BA is still in its nascent stage, we hope that our proposed method will pave the way for further exploration of this difficult optimization problem, and will generate further steps towards initialization-free structure-from-motion.
Appendix
This supplemental material is organized as follows:
Appendix 0.A studies robustness of PoVar with respect to $\eta$ and random initialization.
@@ Appendix 0.B Proof of Lemma 1
The proof in Weber et al. [23] uses the positive-definiteness of $U_{\lambda}$ and $S$ – that still holds, to show that $\mu<1$. It uses the positive-semi-definiteness of $U_{\lambda}^{-\frac{1}{2}}WV_{\lambda}^{-1}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ to conclude that $\mu\geq 0$. For our VarPro formulation, we consider $V_{0}$ instead of $V_{\lambda}$. Nevertheless, it is straightforward that $U_{\lambda}^{-\frac{1}{2}}WV_{0}^{-1}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ is also symmetric positive semi-definite, and then the proof stays almost the same. In details, here is the adapted proof:
@@
On the one hand $U_{\lambda}^{-\frac{1}{2}}WV_{0}^{-1}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ is symmetric positive semi-definite, as $U_{\lambda}$ is symmetric positive definite, and $V_{0}$ is symmetric positive semi-definite. Then its eigenvalues are greater than 00. As $U_{\lambda}^{-\frac{1}{2}}WV_{0}^{-1}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ and $U_{\lambda}^{-1}WV_{0}^{-1}W^{\top}$ are similar,
@@
We extensively use the pOSE formulation [11] for testing our solvers. Recently, the follow-up expOSE formulation [14] has been proposed to override some limitations of pOSE. However, such formulation raises some issues for the scalability analysis. In addition to the fact that the authors do not use the exact VarPro algorithm, expOSE requires an experimental preprocessing step over each image measurements. Without this first step, the exponential function is equal to 00 and the algorithm does not update. Nevertheless, such preprocessing is not feasible, in terms of runtime, when the considered dataset is large enough – which is the topic of our paper, where the number of observations goes up to several tens of millions. Although interesting, expOSE is so far limited to small-scale problems, in line with the problems used by the authors – between $19$ and $30$ poses in their core paper. Extending this pseudo object space error to large-scale formulation is an interesting research direction, orthogonal to our work.
That being said, note that our proposed solver PoVar can be used for solving generic nonlinear problems, and is not restricted to the pOSE formulation.

### v5

@@ 4.3 Power Riemannian manifold optimization
As the projective refinement step in Eq. 7 involves both camera matrices and 3D landmarks in homogeneous forms, we exhibit local scale freedom for both camera and landmark parameters. This necessitates incorporation of the Riemannian manifold optimization framework without which the linearized system of equations is always rank-deficient and unsolvable. While a complete theoretical overview of such framework can be found in [1], we formalize Eq. 7 as:
@@
and then Eq. 20 becomes:
$\begin{pmatrix}\tilde{U}_{\tilde{\lambda}}&&\tilde{W}\\ \tilde{W}^{\top}&&\tilde{V}_{\tilde{\lambda}}\end{pmatrix}\begin{pmatrix}\Delta x_{p}\\ \Delta x_{l}\end{pmatrix}=-\begin{pmatrix}\tilde{b}_{p}\\ \tilde{b}_{l}\end{pmatrix}\,,$ (22)
@@
$\tilde{U}_{\tilde{\lambda}}=\tilde{J}_{p}^{\top}\tilde{J}_{p}+D_{p}^{\top}\tilde{\lambda}_{p}D_{p}\,,$ (23)
$\tilde{V}_{\tilde{\lambda}}=\tilde{J}_{l}^{\top}\tilde{J}_{l}+D_{l}^{\top}\tilde{\lambda}_{l}D_{l}\,,$ (24)
@@
Figure 2: Average performance profiles across all BAL problems for solving the first stage Eq. 5. Given a tolerance $\tau\in\{0.01,0.003,0.001\}$, it represents the percentage of solved problems ($y$-axis) with relative runtime $\alpha$ ($x$-axis). Expansion methods PoVar and PoBA show outstanding speed-accuracy results. Our solver PoVar is competitive, and most notably for the highest accuracy $\tau=0.001$.
Figure 3: Convergence plots of Dubrovnik-253 (left) from BAL datasets with 253 poses and Ladybug-412 with 412 poses, for solving the first stage (Eq. 5). The dotted lines correspond to cost thresholds for tolerance $\tau\in\{0.01,0.003,0.001\}$.
Figure 4: Performance profiles for all BAL problems for solving the first two stages (Eq. 5 and Eq. 7). In each combination, the first solver is used to solve the first stage, and the second solver is used for the second stage. Our proposed combination PoVar followed by a Riemannian expansion method outperforms by a large margin compared to other competitors. Also, our proposed Riemannian-expansion method RiPoBA outperforms the iterative baseline in all cases, given a same solver for the first stage.
Figure 5: Convergence plots of Ladybug-460 (left) from BAL datasets with 460 poses and Venice-1672 with 1672 poses, for solving the second stage (Eq. 7). The dotted lines correspond to cost thresholds for tolerance $\tau\in\{0.01,0.003,0.001\}$. Note that for fair comparison, the initial cost is derived before the first stage. The second cost – that may be higher than the initial one, is the initial cost of the second stage, after the first stage has been run. The runtime includes the time spent to solve the first stage.
@@ 5.1 Implementation
We implement pOSE [^2] We use our custom implementation as the official code of [11] is not publicly available. [11], PoVar and RiPoBA framework in C++, directly on the publicly available implementation of PoBA  [^3] https://github.com/simonwebertum/poba [23]. That leads to fair comparisons with this recent and challenging solver. pOSE differs from PoVar by the use of a direct sparse Cholesky factorization. We also compare to VarPro with the conjugate gradients algorithm, preconditioned with Schur-Jacobi preconditioner, called iterative in our experiments. For the second stage, we compare RiPoBA to the conjugate gradients algorithm preconditioned by Schur-Jacobi preconditioners with Riemannian manifold optimization framework, called RiPCG. Except the solver itself, all implementations share much of the code with [23]. We run experiments on MacOS 14.2.1 with an Intel Core i5 (4 cores at 2GHz).
@@ 5.2 Experimental settings
For each stage, we set the maximum number of iterations to $50$, stopping earlier if a relative function tolerance of $10^{-6}$ is achieved. The damping factor $\lambda$ starts for each stage at $10^{-4}$ and is updated accordingly to the success or failure of the iteration. For expansion methods, we set the maximal order of power series to $20$ and a threshold to $0.01$. For iterative methods, we set the maximum number of inner iterations to $500$. We set $\eta$ for pOSE (Eq. 6) to $0.1$.
Figure 6: Example of dense landmark block in the tangent space with 3 observations from a single landmark. We project the initial dense landmark block (left) onto the tangent space by applying to each $i$-th pose Jacobian $J_{p_{i}}\in\mathbb{R}^{2\times 12}$ the projection $\tilde{x}_{p_{i}}^{\perp}$, and to each $j$-th landmark Jacobian $J_{l_{j}}\in\mathbb{R}^{2\times 4}$ the projection $\tilde{x}_{l_{j}}^{\perp}$. The resulting pose and landmark Jacobians in the tangent space (right) belong to $\mathbb{R}^{2\times 11}$ and $\mathbb{R}^{2\times 3}$ respectively.
@@
As in [23], we leverage the special structure of BA problem and propose a memory-efficient storage. We organize the landmarks into dense blocks. In particular, we apply on each row associated to a landmark the block-matrices of the projection $\tilde{x}^{\perp}_{l_{j}},\tilde{x}^{\perp}_{p_{i}}$ corresponding to the landmark and to the cameras stored in the considered dense landmark block (see Fig. 6).
@@
We extensively evaluate our solver and the baselines on the $97$ real-world bundle adjustment problems from the BAL project page [3]. The number of poses goes from $16$ to $13682$. We refer the reader to Supplemental for further details about these problems. For each problem, we only keep the observation measurements. Pose parameters are randomly drawn from an isotropic Gaussian distribuaton with mean 00 and variance $1$, and landmark parameters are deduced with Eq. 8. Notably and contrary to previous works on initialization-free BA, each solver is ran on the same randomized problem, for fair comparisons.
@@ 5.3 Performance profile
We jointly evaluate both runtime and accuracy with performance profiles [8]. Given a solver, the performance profile maps the relative runtime $\alpha$ to the percentage of problems solved with accuracy $\tau$. Graphically, the performance profile of a given solver is the percentage of problems solved faster than the relative runtime $\alpha$ on the $x$-axis. Let be $S$ and $P$ the sets of solvers and problems, respectively. In practice, we can define the objective threshold for a problem $p$ by:
@@ 5.4 Analysis
Fig. 2 shows the performance profiles for all BAL datasets with tolerances $\tau\in\{0.01,0.003,0.001\}$ to solve Eq. 5. As expected, the direct factorization solver (dashed green) used in Hong et al. [11] shows poor performance. Our solver PoVar (blue) challenges PoBA for the largest tolerance $0.01$, and is by far the most competitive solver for the smallest tolerance $\tau=0.001$, that is for the highest accuracy. For $\tau=0.003$, PoVar slightly outperforms PoBA. We also highlight that our solver outperforms the two competitors associated to the VarPro algorithm, iterative (dashed red) and direct factorization.
Fig. 3 shows two examples of the cost decrease during the first stage. Notably, PoVar (blue) converges much more smoothly than its main challenger PoBA which gets stuck in early iterations. By considering the intersections of the solvers with the cost thresholds (dashed grey lines), it also demonstrates the slowness of the iterative method compared to expansion methods in terms of runtime.
@@
Fig. 4 shows the performance profiles for all BAL datasets with tolerances $\tau\in\{0.01,0.003,0.001\}$ to solve the first two stages, that are Eq. 5 followed by Eq. 7. Note that we compare in this experiment the cost of the second stage only, as the first stage is only used to get an approximated initialization for the projective formulation. As direct factorization shows poor performance during the first stage, we only take into account the most promising combinations of solvers. PoVar followed by RiPoBA (blue) outperforms all the competitors both in terms of runtime and accuracy. The combinations with RiPoBA outperform RiPCG for the highest accuracy $\tau=0.001$ for all relative time greater than $\alpha=2$, that reflects the better convergence of Riemannian expansion method compared to RiPCG. We also note, given a same solver for first stage, RiPoBA outperforms RiPCG during the second stage across all tolerances.
Fig. 5 illustrates on two examples the cost decrease during the second stage. On the left figure (Ladybug-460), the best two solvers in terms of final convergence are built on our framework RiPoBA. On the right figure (Venice-1672), all the solvers using RiPoBA converge to a smaller error than their iterative competitor RiPCG. By considering the intersection between the solvers and the cost thresholds (dashed grey lines), the combination PoVar-RiPoBA outperforms all other combinations, and most notably for the largest problem with $1672$ poses.
@@
The experiments emphasize the high efficacy of our solvers PoVar and RiPoBA, during both first and second stages of the stratified BA problem. Concerning the first stage, the convergence of PoVar is much smoother than its competitors, that explains its larger speed with respect to PoBA to reach the cost thresholds, albeit both are built with power series. Regarding the second stage, for a same given solver in the first stage, our RiPoBA outperforms the preconditioned conjugate gradients with Riemannian manifold optimization framework in terms of speed and accuracy, especially when coupled with PoVar.
@@ 6 Conclusion
We have introduced a novel approach to address the scalability challenge for initialization-free bundle adjustment. Our proposed Power Variable Projection (PoVar) algorithm, theoretically justified, offers new insights to this uncharted problem. By extending recent inverse expansion techniques to the VarPro algorithm on one hand, and to Riemannian manifold optimization on the other hand, we have demonstrated the capability to efficiently solve large-scale stratified BA problem with thousands of cameras. Notably, we achieve state-of-the-art results in terms of speed and accuracy on the real-world BAL dataset. While initialization-free BA is still in its nascent stage, we hope that our method will pave the way for further exploration of this difficult optimization problem, and will generate further steps towards initialization-free structure-from-motion.
6.0.1 Limitations and future work.
First, our analysis applies to the BA problem in which excess outlier point tracks are assumed to have been filtered out as in previous formulations [13, 12, 11, 14]. Second, the 3D reconstruction in Fig. 1 assumes that the intrinsics are known during the metric upgrade stage. In practice, we use the approximated focal lengths given in the BAL dataset, that results in imperfectly accurate illustration. However, some formulations based on pOSE handle unknown intrinsics and could be easily adapted to PoVar.
Acknowledgements
This work was supported in part by the NRF grants funded by the Korea government (MSIT) (No. 2022R1C1C1004907) and in part by the Institute of Information & communications Technology Planning & Evaluation (IITP) under the artificial intelligence semiconductor support program to nurture the best talents (IITP-(2024)-RS-2023-00253914) grant funded by the Korea government(MSIT).
Appendix 0.A studies robustness of PoVar with respect to $\eta$ and random initialization, as well as the scale of the considered problems.
@@ Appendix 0.A Robustness
On the other hand, we link the speed-up of PoVar and RiPoBA, to the size of the problems. Figure 10 represents the performance profiles for solving the first two stages by considering only the BAL datasets with less than $1000$ poses (small to medium-scale problems), and Figure 11 the performance profiles by considering only the BAL datasets with more than $1000$ poses (large-scale problems).
Figure 10: Performance profiles for real-world BAL problems with less than $1000$ poses for solving the first two stages Eq. (6) and Eq. (7).
Figure 11: Performance profiles for real-world BAL problems with more than $1000$ poses for solving the first two stages Eq. (6) and Eq. (7).
@@ Appendix 0.B Proof of Lemma 1
The proof in Weber et al. [23] uses the positive-definiteness of $U_{\lambda}$ and $S$ – that still holds, to show that $\mu<1$. It uses the positive-semi-definiteness of $U_{\lambda}^{-\frac{1}{2}}WV_{\lambda}^{-1}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ to conclude that $\mu\geq 0$. For our VarPro formulation, we consider $V_{0}$ instead of $V_{\lambda}$. Note that the generalized Schur complement is written as: $S=U_{\lambda}-WV_{0}^{{\dagger}}W^{\top}$, where $V_{0}^{{\dagger}}$ is the (Moore-Penrose) pseudo-inverse of $V_{0}$. Nevertheless, it is straightforward that $U_{\lambda}^{-\frac{1}{2}}WV_{0}^{{\dagger}}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ is also symmetric positive semi-definite, and then the proof stays almost the same. In details, here is the adapted proof:
@@
On the one hand $U_{\lambda}^{-\frac{1}{2}}WV_{0}^{{\dagger}}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ is symmetric positive semi-definite, as $U_{\lambda}$ is symmetric positive definite, and $V_{0}$ is symmetric positive semi-definite. Then its eigenvalues are greater than 00. As $U_{\lambda}^{-\frac{1}{2}}WV_{0}^{{\dagger}}W^{\top}U_{\lambda}^{-\frac{1}{2}}$ and $U_{\lambda}^{-1}WV_{0}^{{\dagger}}W^{\top}$ are similar,
@@
We extensively use the pOSE formulation [11] for testing our solvers. Recently, the follow-up expOSE formulation [14] has been proposed to override some limitations of pOSE. However, such formulation raises some issues for the scalability analysis. In addition to the fact that the authors wrongly claim that they use VarPro, expOSE requires an experimental preprocessing step over each image measurements. Without this first step, the exponential function is equal to 00 and the algorithm does not update. Nevertheless, such preprocessing is not feasible, in terms of runtime, when the considered dataset is large enough – which is the topic of our paper, where the number of observations goes up to several tens of millions. Although interesting, expOSE is so far limited to small-scale problems, in line with the problems used by the authors – between $19$ and $30$ poses in their core paper. Extending this pseudo object space error to large-scale formulation is an interesting research direction, orthogonal to our work.
[…]